solve and check system of equations
5) r+s=-6
r-s=-10
10) x+2y=8
x-2y=4

7)c-2d=14
c+3d=9
11)8a+5b=9
2a-5b=-4

Answer by checkley77(12844)

5) r+s=-6
r-s=-10 add
-----------------------
2r=-16
r=-16/2
r=-8 ans.
-8+s=-6
s=-6+8
s=2 ans.
10) x+2y=8
x-2y=4 add
------------------------
2x=12
x=12/2
x=6 ans.
6+2y=8
2y=8-6
2y=2
y=2/2
y=1 ans.

7)c-2d=14
-(c+3d=9) subtract
---------------------------
-5c=5
c=5/-5
c=-1 ans.
-5+2d=14
2d=14+5
2d=19
d=19/2
d=9.5 ans.
11)8a+5b=9
2a-5b=-4 add.
--------------------
10a=5
a=5/10
a=1/2 ans.
8*1/2+5b=9
4+5b=9
5b=9-4
5b=5
b=1 ans.